# Stable-MoE: Lyapunov-based Token Routing for Distributed Mixture-of-Experts Training over Edge Networks

### v1

@@ IV Experimental Results
We evaluate the performance of Stable-MoE in resource-constrained edge networks. The experiments are conducted on two benchmark datasets: SVHN [netzer2011reading] and CIFAR-100 [krizhevsky2009learning], both consisting of $\times32$ colored digit images. The gating network is feedforward layers, and the expert networks are convolutional layers for image classification. We generate the tokens at time slot $t$ by sampling a batch of images from the dataset according to a Poisson distribution with parameter $\lambda$. This sampling strategy ensures balanced label distribution while maintaining the stochastic characteristics of real-time token arrivals. Consider that $J=10$ edge servers with $K=3$, time slot duration $\tau=1\mathrm{s}$, average token arrival rate $\lambda=390\mathrm{tokens/slot}$, effective switched capacitance $\xi_{j}=2\times 10^{-27}$, maximum CPU frequency $f^{\mathrm{max}}=3\mathrm{GHz}$, computational complexity $c_{i}=10^{7}\mathrm{cycles/token}$. We set ${E_{j}^{\mathrm{max}}}\in[3\mathrm{J},15\mathrm{J}]$ and ${E_{j}^{\mathrm{avg}}}\in[1.5\mathrm{J},9.5\mathrm{J}]$, $\forall j\in\mathcal{J}$.
@@
Fig. 2 shows the temporal evolution of token and energy queue backlogs versus $\mathcal{T}$ under Stable-MoE. It is observed that both token queues and energy queues exhibit initial growth followed by stabilization around steady-state values. These results further demonstrate the robust queue stability of Stable-MoE under stochastic token arrivals and heterogeneous energy constraints.
@@
Fig. 3 compares the cumulative system throughput over time of Stable-MoE against four baselines under SVHN and CIFAR-100. It is observed that, Stable-MoE achieves significantly higher cumulative throughput than Strategies A-D, and the performance gain continuously widens over time. For example, the performance gain of Stable-MoE over Strategy D is around $40\%$. This is due to the fact that Strategy D focuses solely on minimizing energy queues without considering the computational workload. As a result, edge servers with smaller energy queues may still be overloaded.
@@ V Conclusion
We have proposed a novel Lyapunov-based token routing framework for distributed MoE systems over resource-heterogeneous edge networks to dynamically optimize token routing and computation frequency allocation. Given stochastic token arrivals, Stable-MoE ensures the long-term stability of both token and energy queues under diverse computational capabilities of distributed edge servers. Experimental results have demonstrated that the proposed approach achieves superior throughput compared with baseline strategies while maintaining queue stability. Overall, this paper serves as a first foray into the dynamic token routing issue of distributed MoE from the perspective of Lyapunov optimization.

### v2

@@ IV Experimental Results
We evaluate the performance of Stable-MoE in resource-constrained edge networks. The experiments are conducted on two benchmark datasets: SVHN [12] and CIFAR-100 [8], both consisting of $\times32$ colored digit images. The gating network is feedforward layers, and the expert networks are convolutional layers. We generate the tokens at time slot $t$ by sampling a batch of images from the dataset according to a Poisson distribution with parameter $\lambda$. This sampling strategy ensures balanced label distribution while maintaining the stochastic characteristics of real-time token arrivals. Consider that $J=10$ edge servers with $K=3$, time slot duration $\tau=1\mathrm{s}$, average token arrival rate $\lambda=390\mathrm{tokens/slot}$, effective switched capacitance $\xi_{j}=2\times 10^{-27}$, maximum CPU frequency $f^{\mathrm{max}}=3\mathrm{GHz}$, computational complexity $c_{i}=10^{7}\mathrm{cycles/token}$. We set ${E_{j}^{\mathrm{max}}}\in[3\mathrm{J},15\mathrm{J}]$ and ${E_{j}^{\mathrm{avg}}}\in[1.5\mathrm{J},9.5\mathrm{J}]$, $\forall j\in\mathcal{J}$. These non-uniform energy budgets intrinsically constrain $f_{j}(t)$ and correspondingly limit $d_{j}^{\text{com}}(t)$ of each expert, thereby leading to the heterogeneous computational capabilities of edge servers.
@@
Fig. 2 shows the temporal evolution of token and energy queue backlogs versus $\mathcal{T}$ under Stable-MoE. It is observed that both token and energy queues exhibit initial growth followed by stabilization around steady-state values. These results further demonstrate the queue stability of Stable-MoE under stochastic token arrivals and heterogeneous energy constraints.
Fig. 3 compares the cumulative system throughput over time of Stable-MoE against four baselines under SVHN and CIFAR-100. It is observed that, Stable-MoE achieves significantly higher cumulative throughput than Strategies A-D, and the performance gain continuously widens over time. This is due to the fact that Strategy D focuses solely on minimizing energy queues without considering the computational workload. As a result, edge servers with smaller energy queues may still be overloaded.
@@ V Conclusion
We have proposed a novel Lyapunov-based token routing framework for distributed MoE systems over resource-heterogeneous edge networks to dynamically optimize token routing and computation frequency allocation. Given stochastic token arrivals, Stable-MoE ensures the long-term stability of both token and energy queues under diverse computational capabilities of distributed edge servers. Experimental results have demonstrated that Stable-MoE achieves superior throughput compared with baseline strategies while maintaining queue stability. Overall, this paper serves as a first foray into the dynamic token routing issue of resource-heterogeneous MoE from the perspective of Lyapunov optimization. Going forward, future directions include the designs of dynamic routing strategies under sophisticated model architecture and over resource-limited wireless networks.